This is from a mom trying to help child with algebra - the question:
Solve the system algebraically. Identify the system and give the number of solutions. (2x-2y=-4
(6x-6y=-24
Is it consistent/dependent 2 solutions
consistent/independent 1 solution
inconsistent 0 solutions
inconsistent 2 solutions
Please help

Answer by stanbon(75887)

Identify the system and give the number of solutions.
(2x-2y=-4
(6x-6y=-24
Is it consistent/dependent 2 solutions
consistent/independent 1 solution
inconsistent 0 solutions
inconsistent 2 solutions
----------------
1st: 2x-2y=-4
Divide thru by 2 to get:
x-y=-2
2nd: 6x-6y=-24
Divide thru by 6 to get:
x-y=-4
-------------
These equations represent two lines that are parallel.
No x,y pairs satisfy both equations.
There are no solutions to the system.
The system is inconsistent.
Cheers,
Stan H.